# Frontiers in Probabilistic Inference: Sampling Meets Learning

### v1

@@ Appendix A Proof for Equation 7
$\frac{s^{N}}{(2\pi)^{N}h^{n}C}\int_{V}d\vec{q}e^{-\frac{U(\vec{q})}{k_{B}T}}\delta^{N}\Big(\vec{Q}-f(\vec{q})\Big)\sqrt{\frac{(k_{B}T)^{n}(2\pi)^{n}}{\text{det}M^{-1}(\vec{q})}}\sqrt{\frac{(k_{B}T)^{-N}(2\pi)^{N}}{\text{det}\big(J_{f}(\vec{q})M^{-1}(\vec{q})J_{f}^{\top}(\vec{q})\big)}}\\ e^{-\frac{1}{k_{B}T}P^{\top}(R^{-1}(\vec{Q}))^{\top}\big(J_{f}(\vec{q})M^{-1}(\vec{q})J_{f}^{\top}(\vec{q})\big)^{-1}R^{-1}(\vec{Q})P}$ (13)
@@ Appendix B Coarse Grain Function Candidates
B.1 Isometric Invariance
We assume that the standard case of $M$ is a position-independent diagonal matrix, of the same form as 2. Define the set of isometries that also commute with $M$ to be $G$ be all functions of $g(\vec{q})=R\vec{q}+b$ where $RMR^{\top}=M$ Here we will show that if $f$ is transitive on level sets, such that for every. If $f$ is transitive on level sets, so that $\forall\vec{q}_{1}\ \forall\vec{q}_{2}\ \ f(\vec{q}_{1})=f(\vec{q}_{2})\rightarrow\exists(g\in G)\ s.t.\ \vec{q_{1}}=g(\vec{q_{2}})$, and $f$ is invariant on the subset of $G$ used within level sets, so that $f=f\circ g$, then equation 6 can be satisfied. We can show that $J_{f}(\vec{q}_{1})MJ^{\top}_{f}(\vec{q}_{1})=J_{f}(\vec{q}_{2})MJ^{\top}_{f}(\vec{q}_{2})$ for any microstates $\vec{q}_{1},\vec{q}_{2}\in f^{-1}(\vec{Q})$ in the preimage of the same macrostate as follows:
@@ B.1 Isometric Invariance
and the Jacobian of function composition is each Jacobian composed.
@@
This proof shows the existence of an $R(Q)$ for any choices of $f$ that satisfies the conditions. ∎
@@ B.2 Distance Latent Space
Consider a coarse grain function $f(\vec{q})$ that takes the pairs of distances between pairs of atoms. $f$ is isometric invariant almost by definition, so an $R$ should exist. We can find it by
@@
Note that if $ab\cap cd$ do not share endpoints, then its just zero since one of the derivatives will always be zero. If they do, then let $b$ be the atom that they share.
@@
Thus if $f$ provides enough edges so that the structure is "rigid", then the angles can be determined from the cosine rule and a formula for $R$ is trivial.

### v2

@@ Appendix A Proof for Equation 8
$\frac{s^{N}}{(2\pi)^{N}h^{n}C}\int d\vec{q}e^{-\frac{U(\vec{q})}{k_{B}T}}\delta^{N}\Big(\vec{Q}-f(\vec{q})\Big)\sqrt{\frac{(k_{B}T)^{n}(2\pi)^{n}}{\text{det}M^{-1}(\vec{q})}}\sqrt{\frac{(k_{B}T)^{-N}(2\pi)^{N}}{\text{det}\big(J_{f}(\vec{q})M^{-1}(\vec{q})J_{f}^{\top}(\vec{q})\big)}}\\ e^{-\frac{1}{k_{B}T}\frac{1}{2}P^{\top}(R^{-1}(\vec{Q}))^{\top}\big(J_{f}(\vec{q})M^{-1}(\vec{q})J_{f}^{\top}(\vec{q})\big)^{-1}R^{-1}(\vec{Q})P}$ (15)
@@ Appendix B Coarse Grain Function Candidates
B.1 Isometric and Maximal Invariance
Define the set of isometries $G$ that also commute with $M$ to be all functions of $g(\vec{q})=R\vec{q}+b$ where $RMR^{\top}=M$. If $f$ is transitive on level sets, so that $\forall\vec{q}_{1}\ \forall\vec{q}_{2}\ \ f(\vec{q}_{1})=f(\vec{q}_{2})\rightarrow\exists(g\in G)\ s.t.\ \vec{q_{1}}=g(\vec{q_{2}})$, and $f$ is invariant on the subset of $G$ used within level sets, so that $f=f\circ g$, then equation 7 can be satisfied. We can show that $J_{f}(\vec{q}_{1})MJ^{\top}_{f}(\vec{q}_{1})=J_{f}(\vec{q}_{2})MJ^{\top}_{f}(\vec{q}_{2})$ for any microstates $\vec{q}_{1},\vec{q}_{2}\in f^{-1}(\vec{Q})$ in the preimage of the same macrostate as follows:
@@ B.1 Isometric and Maximal Invariance
and the Jacobian of function composition is the composition of the Jacobians.
@@
This proof shows the existence of an $R(Q)$ for any choice of $f$ that satisfies the conditions. ∎
@@ B.2 Distance Latent Space
Consider a coarse-grained function $f(\vec{q})$ that takes the distances between pairs of atoms. $f$ is isometric invariant almost by definition, so an $R$ should exist. We can find it by defining
@@
Note that if $ab\cap cd$ do not share endpoints, then it is just zero since one of the derivatives will always be zero. If they do, then let $b$ be the atom that they share.
@@
Thus, if $f$ provides enough edges so that the structure is "rigid", then the angles can be determined from the cosine rule, and a formula for $R$ is trivial.
B.2.1 Distance Latent Space Results
We trained a simple feed-forward network on the 1st, 6th, and 10th Carbon-alpha atom of Chignolin (1UAO) at 300K. The input was two distances of the sixth atom to the other two, and the cosine of the angle. The coarse-grain function is:
$f(\vec{q})=\begin{bmatrix}\lVert\vec{q}_{1}-\vec{q}_{6}\rVert\\ \lVert\vec{q}_{10}-\vec{q}_{6}\rVert\\ \frac{(\vec{q}_{1}-\vec{q}_{6})\cdot(\vec{q}_{10}-\vec{q}_{6})}{\lVert\vec{q}_{1}-\vec{q}_{6}\rVert\lVert\vec{q}_{10}-\vec{q}_{6}\rVert}\end{bmatrix}$
The model outputs a single scalar, free energy. The gradient of the free energy was then trained using force matching Noid2008.
$L(\theta)=\frac{1}{M}\sum_{i=1}^{M}\Big\lVert\nabla_{Q}W(f(D_{i}),\theta)-\nabla_{Q}F(f(D_{i}))\Big\rVert^{2}$
where $\nabla_{Q}F$ is the expression in Eq. 10.
Figure 1: Ground truth and model prediction histogram of the probabilities in latent space. Units of length are in nanometers. Red points are the strided positions of the ground truth data. The unnormalized model probability was evaluated using the Boltzmann distribution.